OpenStudy (anonymous):

3x+y=20 y=2x substitute

OpenStudy (anonymous):

5x=20 x=20/5

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Because it is substitution, you're putting the y value of one equation in as the y value in another. So you'd be putting the y from y=2x into the y from 3x+y=20. 3x+(2x)=2= 5x=20 divide by 5 x=4. Then you'd put that back into the y=2x y=2(4) y=8. So those two equations intersect at (4,8)